### v1

# GUICourse: From General Vision Language Model to Versatile GUI Agent

Abstract
Utilizing Graphic User Interface (GUI) for human-computer interaction is essential for accessing a wide range of digital tools. Recent advancements in Vision Language Models (VLMs) highlight the compelling potential to develop versatile agents to help humans finish GUI navigation tasks. However, current VLMs are challenged in terms of fundamental abilities (OCR and grounding) and GUI knowledge (the functions and control methods of GUI elements), preventing them from becoming practical GUI agents. To solve these challenges, we contribute GUICourse, a suite of datasets to train visual-based GUI agents from general VLMs. First, we introduce the GUIEnv dataset to strengthen the OCR and grounding capabilities of VLMs. Then, we introduce the GUIAct and GUIChat datasets to enrich their knowledge of GUI components and interactions. Experiments demonstrate that our GUI agents have better performance on common GUI tasks than their baseline VLMs. Even the small-size GUI agent (with 3.1B parameters) can still work well on single-step and multi-step GUI tasks. Finally, we analyze the different varieties in the training stage of this agent by ablation study. Our source codes and datasets are released at https://github.com/yiye3/GUICourse.
1 Introduction
Graphical User Interfaces (GUIs) serve as a pivotal medium for facilitating human-computer interactions, playing a pivotal role across diverse applications. Different GUI systems (e.g. websites and smartphones) usually use consistent icons and operation logic, which makes users ignore the complexity of the underlying processes and promotes skill transfer between different systems. GUI agents, which are constructed for GUI navigation tasks, not only liberate humans from tedious and repetitive operations, but can also be regarded as an important method for using GUI tools, thereby expanding the capabilities of large language models (LLMs).
Recent works have explored training GUI agents in various scenarios, such as computer control [14], web browsing [7, 44], and smartphone navigation [31, 36, 39, 43]. These agents can be classified based on their input modalities into textual-based (e.g., HTML, source code) GUI agents and visual-based (e.g., screenshots) GUI agents. We argue that visual-based GUI agents have two advantages over textual-based methods: (1) Ease of acquisition: Obtaining screenshots of GUI systems is straightforward, whereas accessing structured text, such as the code behind the iOS system, is challenging. (2) Improved transferability: GUIs typically utilize similar visual elements to denote similar functions across different systems.
The performance of vision-based agents heavily relies on the fundamental capabilities and internal knowledge of current Vision Language Models (VLMs). Although VLMs [20, 26, 52, 17, 13] demonstrate impressive capabilities in various tasks such as image captioning [18] and visual question answering [21], these capabilities may not suffice for developing efficient visual-based GUI agents. For example, we investigate Qwen-VL-Chat, a powerful open-sourced VLM, on its ability to finalize GUI navigation instructions. As illustrated in Figure 1, it shows the two challenges of current VLMs: (1) Their OCR and grounding abilities are not supported to locate the website elements or designated text accurately; (2) They lack a comprehensive understanding of the functions and control mechanisms of website widgets.
Figure 1: An example of several GUI tasks. Qwen-VL-Chat gives an incorrect bounding box for "Log in", which means its OCR and grounding capabilities are insufficient. Also, it can’t finish GUI navigation tasks such as "find the products" and "accept the cookies" because it lacks the knowledge of GUI control actions. Trained on our datasets, our GUI agent Qwen-GUI performs these tasks well.
In this work, We contribute GUICourse, a group of complete datasets to help improve VLMs’ fundamental abilities and GUI knowledge. GUICourse is composed of three datasets: (1) GUIEnv, a large-scale dataset for improving VLMs’ OCR and grounding abilities, including 10M website page-annotation pairs as pre-training data and 0.7M region-text QA pairs as SFT data; (2) GUIAct, a GUI navigation dataset in website and android scenarios for enhancing VLMs’ knowledge of GUI systems, including 67k single-step and 15k multi-step action instructions. (3) GUIChat, a conversational dataset for improving the interaction skills of GUI agents, including 44k single-turn QA pairs and 6k multi-turn dialogues with text-rich images and bounding boxes. Based on the proposed dataset, we train several GUI agents based on different VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V). We demonstrate that our GUI agents have better performance on Mind2Web and AITW than their baseline VLMs. Then, we conduct ablation studies on our GUI agent MiniCPM-GUI, comparing the impact of different GUIEnv pre-training data amounts, image resolutions, and mixed GUIChat data. We demonstrate that GUIEnv data is effective in improving VLMs’ OCR and grounding abilities. These abilities, as well as high resolution, significantly increase the performance of GUI tasks.
2 Related Works
2.1 Vision Language Models
Currently, most Vision-Language Models (VLMs) use a bridge module to connect a vision encoder, such as CLIP-ViT [8, 28], and a language model, such as Vicuna [38] and Qwen [3]. The bridge module maps the visual features from the vision encoder into the embedding space of the large language model with an MLP layer, exemplified by LLaVA [21], or an attention-based resampler, such as BLIP-2 [18] and Qwen-VL [4]. Uniquely, Fuyu-8B [5] removes the external vision encoder and uses a decoder-only transformer architecture with a unified space for textual tokens and pixel information, enabling Fuyu-8B to process images of any resolution. Based on this, subsequent research efforts try to enhance VLMs in various domains, including visual detail recognition [40, 49], reliable responses [37, 47], object grounding [26], multilingual capabilities [13] and model efficiency [53]. However, many VLMs still have shortcomings when (1) recognize small-size text in high-resolution images, and (2) provide the accurate location of the designated area at the pixel level. In this work, we provide the GUIEnv dataset to improve the VLMs’ OCR and grounding abilities with high-resolution website screenshots.
2.2 GUI Agents
GUI agents are expected to help humans finish tasks on different GUI systems, such as simplified websites [33, 19], simulated environments [44, 51], or real-world websites and smartphones [7, 31]. Nowadays, the GUI agents are developing quickly due to the Large Models (LMs), Big Data, and data-driven methods. Some agents [10, 34, 23, 15, 41, 50, 43] using the train-free method and depending on closed-sourced LMs (e.g., GPT4 and GPT-4V [25]) to achieve GUI navigation with strategy design and prompt engineering. Also, some agents [14, 24, 27] finish GUI tasks based on small-size open-sourced models, using data-driven methods (e.g., supervised fine-tuning).
Perception and control are the two important points for GUI agents. According to the perception modalities of environments, the GUI agents can split into text-only [11, 7], image-text combined [14, 9], and vision-only [32, 48, 12] methods. As for the control, GUI agents can split into position-free and position-needed methods. The position-free methods [7, 44] generally provide indexes for every element in the GUI systems, and agents can finish actions by multiple choices. While the position-needed methods [32, 48, 12] mean agents need to give the position information to finish some actions, such as "click" and "tap". In this work, we aim to train GUI agents with vision-only inputs and position-needed actions using data-driven methods.
To train a visual-based GUI agent using data-driven methods, lots of GUI navigation data with visual modal information (e.g., screenshots) are needed. Some works provide GUI navigation datasets including visual modal information, but their datasets usually have one or more of the following problems: (1) The GUI environments are too easy, and far from real-world scenarios [19, 14]. (2) The datasets are focused on a narrow domain or scenario [44, 35, 31]. (3) The size of datasets is too small to train GUI agents [7]. In this work, we contribute the GUIAct dataset to solve these three problems. Besides that, we also provide the GUIChat dataset to improve the GUI agents’ natural language interaction abilities.
3 GUICourse Construction
In this section, we introduce GUICourse, a group of datasets for improving VLMs’ OCR and grounding abilities, enhancing VLMs’ GUI system knowledge, and helping GUI agents’ interaction. As shown in Table 1, our GUICourse is the first large-scale corpus for the entire pipeline to train GUI agents from general VLMs, which includes GUIEnv, GUIAct, and GUIChat.
Table 1: Statistics of the GUICourse compared with existing datasets. The columns indicate: whether data includes generating and understanding the content in designated regions and full pages (Region and Page), the number of training samples (Num.), whether real-world scenarios are used (Real.), the number of instructions in website and android scenarios (Web Inst. and And. Inst.), whether there are text-rich images and bounding boxes in dialogues (Text-Rich and w/ Bbox).
Dataset OCR & Grounding GUI Knowledge Interaction
Region Page Num. Real. Web Inst. And. Inst. Text-Rich w/ Bbox
MiniWoB++ [19] - 100 -
WebShop [44] - 12k -
META-GUI [36] - - 1,125
MIND2WEB [7] - 2,350 -
AITW [31] - - 715k
WEBLINX [22] - 2,337 -
SeeClick [6] <1M - - -
ScreenAI [2] - - - -
Ferret-UI [46] <1M - - -
GUICourse (Ours) 10.7M 73k 9,157
Figure 2: Some examples to show the data formats of our GUIEnv datasets. The GUIEnv-global task only provides a screenshot website, and the models need to predict all the texts and their positions according to the layout sequence. The GUIEnv-local task requires models to predict the region (a box represented by a left-top point and a right-down point) given the designated text or predict the text given the designated box.
3.1 GUIEnv: A Large Scale Dataset for Fundamental Abilities
OCR and grounding are two fundamental abilities for visual-based GUI agents’ perception and control. To augment these abilities, we create a large-scale dataset called GUIEnv with about 10M pre-training data GUIEnv-global and 0.7M SFT data GUIEnv-local. As shown in Figure 2, each sample in GUIEnv-global is a long text with all the describable content on the full page, including text, grounding information, and layout sequence. Each sample in GUIEnv-local is a QA pair on a designed region as the "text2bbox" or "bbox2text" task.
Data Collection.
We collected 4M URLs from the C4 [29] and employed Playwright [^1] https://github.com/microsoft/playwright for rendering, ultimately producing 10M annotated screenshots[16]. The 10M annotated screenshots correspond to the 10M samples of the GUIEnv-global data. As for the GUIEnv-local data, we select about 50k website screenshots with their elements from C4. These screenshots are usually huge due to they include the complete content of the websites. We process the data in three steps: (1) crop them to smaller partitions less than $1920\times 1080$ resolution. (2) remove images with elements less than 10. (3) randomly simply 10 elements with texts and positions in each screenshot as the data of "text2bbox" and "bbox2text" tasks. Finally, we acquired 0.7M instructions for the GUIEnv-local data.
3.2 GUIAct: A Multi-Scenario Dataset for GUI Knowledge
GUI systems’ knowledge, which includes the functions and the control methods of GUI elements, is important for GUI agents. Aiming at bridging the gap between VLMs and GUI agents, we contribute the GUIAct dataset to enhance GUI knowledge. The GUIAct dataset includes GUI navigation tasks in two scenarios (website and smartphone), which can be split into three partitions: "web-single", "web-multi", and "smartphone". GUI navigation tasks require agents to interact with GUI environments using actions, according to a task instruction, as shown in Figure 3. In this work, we only consider the GUI agents to interact with image-only environments. We only use the screenshots of GUI systems in the viewport, without any unseen elements, which use the same environments as humans do. Different GUI systems generally have diverse actions (e.g. "click" is common on websites while "tap" is common on smartphones). We define a unified action space for different GUI systems, including 11 types of actions, and the details are shown in Appendix A.1.
Figure 3: An example of our GUIAct dataset with 17 steps of action-screenshot pairs. Each step corresponds with one website screenshot and one or more actions. Some actions (e.g., “click”) have their location information, represented by a bounding box.
Data Collection.
We contribute the GUIAct (web-single) dataset in four steps: (1) Websites selection. We use GPT-4 to gather diverse scenarios (e.g., online shopping and E-learning) and URLs of using the internet. Then, we use these URLs as seeds to expand new websites by their hyperlinks. We acquire 50 domains and 13k websites in this step. (2) Captures acquisition. We use web snapshot tools to respond to website HTML, interactive elements, and screenshots based on their URLs. (3) LLM-Auto Annotation. We use GPT-4V to obtain single-step instruction-action pairs for each website. We give two images to GPT-4V in each request: an origin screenshot and a revised screenshot with interactive element identification. We get about 70k instruction-action pairs based on the 13k website screenshots. (4) Data checking. We hire annotators to check the automatic instruction-action pairs’ quality. The annotators are requested to revise the inaccurate items or abandon them if they are hard to revise. After that, the accuracy of GUIAct (web-single) data improves from 55% to 92%, based on our sampled results. Finally, we get around 67k single-step action instructions in website scenarios, corresponding to 67k training samples.
We contribute the GUIAct (web-multi) dataset in three steps: (1) Websites selection. We concept 8 top-level web scenarios for information searching: Shopping, Diet, Stay, Travel, Education, Health, Job, and Entertainment. Based on the 8 top-level scenarios, we collect 32 sub-scenarios and 121 well-known websites. (2) Questions acquisition. We use LLMs (GPT-3.5 and Claude2) to generate instructions for each website. We collect 8k high-level instructions for searching special information on the corresponding websites. To finish each instruction needs multi-turn interaction with the website. (3) Crowd-sourced Annotation. We develop an annotation tool as a web browser plugin. Then, we hire annotators to execute operators to finish the instructions on the corresponding websites. If the instructions are inaccurate or have no answer, the annotators will revise or abandon them. Finally, We get 5,696 multi-step action instructions in website scenarios, corresponding to 44k training samples.
We contribute the GUIAct (smartphone) dataset using a subset of the AITW [31] dataset. We select the data with the "General" tag and filter out the screenshots without the bottom navigation bar (because we will convert the "go back" and "go home" actions to "tap"). Then, we convert the origin actions to our action space. Finally, we get 9,157 multi-step action instructions in smartphone scenarios, corresponding to 67k training samples.
3.3 GUIChat: A Text-Rich Multi-Modal Dataset for Interaction
Natural language interaction is important for better using GUI agents. It’s necessary to mix conversation data during the training stage of GUI agents for better interaction. We introduce the GUIChat dataset, which includes lots of text-rich website screenshots and grounding boxes in dialogues. This dataset has about 44k single-turn QA pairs and 6k multi-turn dialogues in four aspects: visual information queries, human-centric issues, world knowledge inquiries, and complex reasoning tasks.
Data Collection.
We contribute this dataset in three steps: (1) Web Page Screenshot Acquisition. Utilizing Playwright to render web pages and obtain screenshots, same as the collection process of GUIEnv. (2) Textual Representation Extraction. Extracting essential structured information and text with coordinates from a redundant DOM tree. (3) Question-Answer Pair Generation. Leveraging GPT-4 to construct question-answer pairs from the textual representations of the current website. The textual representations include the elements in the website with the position information. We use GPT-4 rather than GPT-4V to achieve this step because GPT-4 can get higher accuracy than GPT-4V when given accurate textual representations.
4 Experiments
We train several GUI agents with our GUICourse (GUIEnv, GUIAct, and GUIChat datasets) based on three VLMs: Fuyu-8B [5], Qwen-VL [3], and MiniCPM-V [1]. First, we explore the performance of our GUI agents for common GUI tasks: Mind2Web [7] and AITW [6]. Then, we do ablation studies of the whole training process based on MiniCPM-V, analyzing the influence of various factors. Finally, we use two examples to show the generalization of our GUI agents.
4.1 Performance on GUI Navigation Tasks
We finetune our GUI agents (Qwen-GUI, Fuyu-GUI, and MiniCPM-GUI) and their baseline VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V) on the train set of common GUI tasks (Mind2Web and AITW) respectively and then evaluate their performance. We mainly compared our GUI agents with their baseline VLMs. Meanwhile, we add some previous works for a comprehensive comparison: We compare with MindAct [7], SeeClick [6], and GPT-4 [25] in the Mind2Web task; We compare with SeeClick, ChatGPT-CoT [48], and GPT-4V [42] in the AITW task.
Evaluation Datasets and Metrics.
Mind2Web and AITW are two commonly used GUI tasks. Mind2Web has 2000 tasks with high-level instructions in website scenarios. AITW has 30k instructions and corresponding 715k operation trajectories in smartphone scenarios. To keep consistency with previous works, we choose the processed version of these two tasks by SeeClick [6] and use the same evaluation methods as it does. We remove the "General" column of AITW in the evaluation because we convert and use a subset of the AITW-General dataset into our GUIAct dataset.
Table 2: Results of our GUI agents on the Mind2Web datasets.
Models w/o Cross-Task Cross-Website Cross-Domain
HTML Ele.Acc Op.F1 StepSR Ele.Acc Op.F1 StepSR Ele.Acc Op.F1 StepSR
MindAct 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
GPT-4 41.6 60.6 36.2 35.8 51.1 30.1 37.1 46.5 26.4
SeeClick 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
Qwen-VL 23.2 88.4 20.3 16.8 85.8 14.0 14.1 85.2 12.3
Qwen-GUI 27.9 88.1 24.4 19.3 84.1 15.6 20.5 85.4 17.5
Fuyu-8B 8.3 83.9   6.6   4.8 81.3   4.0   3.6 83.0   3.0
Fuyu-GUI 19.1 86.1 15.6 13.9 80.7 12.2 14.2 83.1 11.7
MiniCPM-V 11.0 85.6   8.5   8.2 78.2   6.0   6.5 81.4   5.2
MiniCPM-GUI 23.8 86.8 20.8 20.3 81.7 17.3 17.9 74.5 14.6
Table 3: Results of our GUI agents on the AITW datasets.
Models Modality Install Google Apps Single WebShopping Overall
ChatGPT-CoT Text   4.4 10.5   9.4   8.4   8.2
GPT-4V Image 42.6 49.8 72.8 45.7 52.7
SeeClick Image 66.4 54.9 63.5 57.6 60.6
Qwen-VL Image 70.4 57.8 70.1 64.7 65.8
Qwen-GUI Image 70.3 61.2 71.6 66.1 67.3
Fuyu-8B Image 45.9 40.0 47.2 40.8 43.5
Fuyu-GUI Image 50.9 41.6 45.7 43.8 45.5
MiniCPM-V Image 50.2 45.1 56.2 44.0 48.9
MiniCPM-GUI Image 62.3 46.5 67.3 57.5 58.4
Results.
The results in Table 2 and Table 3 demonstrate that GUICourse helps GUI agents achieve better performance than their baseline VLMs: (1) In the Mind2Web task, Qwen-GUI exhibits a remarkable increase of $2$-$5$ points in step success rate (StepSR) compared to its baseline Qwen-VL. Fuyu-8B and MiniCPM-V, which have weaker fundamental abilities, show more improvements of approximately $10$ points. The HTML-based methods (such as MindAct) achieve higher Ele.Acc and StepSR scores because they simplify the problem and use multiple-choice QA formulation. Conversely, the methods without HTML need to predict the position of the target elements based on the provided screenshots. (2) In the AITW task, our agents still have better performance than their baseline VLMs. While, these improvements are relatively smaller compared to the Mind2Web task, likely due to the fact that the samples in AITW are comparatively easier.
4.2 Ablation Study
We show two ablation experiments in this section. First, we adjust the amount of the GUIEnv-global dataset and analyze the final performances of MiniCPM-GUI. Secondly, we use the general VLM MiniCPM-V as a baseline, adding the GUIAct dataset, improving the image resolution, and mixing the GUIChat dataset successively. We analyze the influences of different datasets and factors.
Evaluation Datasets and Metrics.
We randomly sample 1.8k samples from the GUIEnv-local dataset as the test data of Bbox2Text and Text2Bbox tasks, testing the OCR and grounding ability of agents. We evaluate the Bbox2Text task by EM and F1 scores (the same as SQuAD MRC [30] task) and evaluate the Text2Bbox task by IoU scores (IoU@0.2, 0.5, and 0.7). We split about 1.4k, 1k, and 2k samples from the "web-single", "web-multi", and "smartphone" partitions in GUIAct as test data. We use the type exact match score (Type EM), step success rate (StepSR), and click accuracy (Cli. Acc) as metrics. Type EM is the accuracy of the actions’ names. StepSR is the average score of all the actions’ success rates. Cli.Acc is the average score of the success rate of the "click" and "tap" actions, which reflects the ability of grounding.
Table 4: The performance of MiniCPM-GUI with different amounts of the GUIEnv data.
Amount Bbox2Text Text2Bbox Web-Single
EM Score F1 Score IoU@0.2 IoU@0.5 IoU@0.7 Type EM Cli.Acc StepSR
w/o GUIEnv 14.92 30.49   9.87   2.15   0.32 92.70 49.80 52.84
w/ 2.5M GUIEnv 35.87 57.38 51.39 25.32   9.76 91.84 66.40 64.82
w/ 5.0M GUIEnv 40.21 62.61 60.73 35.09 16.09 92.98 71.20 68.30
w/ 7.5M GUIEnv 43.60 63.46 65.34 40.67 20.06 91.56 71.50 68.58
w/ 10M GUIEnv 44.12 64.78 68.02 47.96 23.28 91.77 74.90 70.57
Table 5: The performance of MiniCPM-GUI with different variables.
Variables Web-Single Web-Multi Smartphone Mean
Type EM Cli.Acc StepSR Type EM Cli.Acc StepSR Type EM Cli.Acc StepSR StepSR
+ GUIAct 89.1 55.2 52.4 62.3 48.6 42.0 64.5 43.3 29.6 41.3
+ High Resolution 91.8 69.2 72.9 68.2 47.1 42.4 72.6 52.0 47.5 54.3
+ GUIChat 91.8 70.6 74.9 67.0 47.5 45.4 71.7 53.3 44.7 55.0
Results.
The performances of MiniCPM-GUI with different settings are shown in Table 4 and Table 5. We get three findings from these results:
(1) Our GUIEnv dataset is useful to improve both the fundamental abilities of VLMs and the GUI navigation ability. As shown in Table 4, the scores of the "text2bbox" and "bbox2text" tasks, as well as the performance on GUI navigation tasks (web-single) is rising during the increase in the amount of GUIEnv data. We can see that the improvement is huge from "without GUIEnv" to "2.5M GUIEnv", which demonstrates that OCR and grounding abilities are necessary. The amount of GUIEnv data does not influence the "Type EM" because predicting action names is position-free.
(2) High resolution is important for GUI navigation tasks. The default resolution of MiniCPM-V is $448\times 448$, and it can be improved to $1344\times 1344$ by the flexible patching and slicing strategy [45]. As shown in Table 5, there are huge improvements in the "web-single" and "smartphone" tasks after improving the image resolution. We also find that the “web-manual" task has little improvement. That may be because it has more complex instructions and environments, which only improve the resolution is not enough.
(3) Our GUIChat dataset is helpful for tasks in website scenarios, although its main goal is to improve the interaction ability of GUI agents. As shown in Table 5, we mix GUIChat data in the SFT state of GUI agents, and the "StepSR" improved by $2$-$3$ points on "wen-single" and "web-multi" tasks.
4.3 Case Study
Figure 4: An example of a multi-turn conversation with our GUI agent, MiniCPM-GUI. We first ask the agent about possible operations on this website, then let it explain the terminologies in the screenshot. After that, we test the grounding ability of the agent by giving it a GUI navigation instruction and an information search question.
Multi-Turn Dialogue.
As shown in Figure 4, we show a multi-turn dialogue example of our GUI agents. This example uses several unseen website screenshots and illustrates its proficiency across four key areas: (1) Analyzing the functions of GUI systems’ elements; (2) Answering questions using its internal knowledge; (3) Creating actionable steps to accomplish specific tasks; (4) Extracting textual information from the screenshot along with its coordinates through grounding.
Figure 5: An example of our GUI agent executing the same instruction with different environments. The numbers in the "<point></point>" are the related positions from 0 to 1000, and the "(500,500)" corresponds to the center point of the screenshot.
Simulated Environment.
To evaluate the robustness of our GUI agents, we use a simulated smartphone environment by Android Studio [^2] https://developer.android.com/ to do interactive testing. As shown in Figure 5, we let our GUI agent execute the same instruction with different initial environments. The execution processes are shown in the the left three and the right two screenshots respectively. When there is no "setting" icon on the initial environment, our GUI agent swipes first to find the "setting" icon. While the "setting" icon appears on the initial environment, our GUI agent taps directly. That means our agents can distinguish the detailed changes and know the icon of "settings", not only learn the fixed action sequences in the training data. A long-chain example is shown in Appendix B.3.
5 Conclusion
In this work, we construct GUICourse, a group of complete datasets (GUIEnv, GUIAct, and GUIChat) to help train GUI agents from general VLMs from three aspects respectively: improving the OCR and grounding abilities of VLMs, enhancing VLMs’ GUI knowledge, and improving GUI agents’ interaction ability. We train several GUI agents (Qwen-GUI, Fuyu-GUI, and MiniCPM-GUI) with different VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V). We evaluate the GUI agents on common GUI tasks (Mind2Web and AITW), conduct ablation studies on the self-partitioned test sets, and try them on simulated Android environments. Our experiments demonstrated that our datasets are helpful for better GUI navigation ability.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Appendix C.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Appendix C.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://github.com/yiye3/GUICourse.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Appendix B.1.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] The training cost of large models is too high, which makes it’s difficult to run experiments multiple times.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Appendix B.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Section 2, Table 1
    2. (b)
      Did you mention the license of the assets? [No]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] https://github.com/yiye3/GUICourse.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Appendix C.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] Appendix C.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] Appendix A.3.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Appendix A.3.
Appendix A Datasets
A.1 Action Space
In previous works, various definitions of the GUI agents’ action space have been proposed. However, these approaches often suffer from one or more of the following issues: (1) The partitioning of actions is disorganized and lacks coherence. (2) The action space fails to encompass all possible scenarios in GUI systems. To address these limitations, our work introduces a unified action space for GUI systems. Specifically, we have designed seven distinct groups of actions tailored for task-oriented GUI agents, as depicted in Figure 6:
(1) Pointing to something (e.g., buttons, icons): This action allows the agent to shift focus to specific elements within the GUI system, such as clicking on a search box or selecting empty spaces on websites. These actions include "click", "hover", and "tap".
(2) Inputting something: Agents can input relevant information into the GUI system, such as inputting search queries in a search box. These actions include "input", "type", and "paste".
(3) Browsing through pages: These actions enable the agent to navigate through pages within the GUI system, including page-up and page-down functionalities. These actions include "scroll", "drag", and "swipe".
(4) Logging information: Agents can log key information that aids in accomplishing the task, such as writing texts or copying essential paragraphs. These actions include "select text", "copy", and "write (by OCR)".
(5) Submitting (tables, search text): Agents can submit tables or the search context within the GUI system. We use the "enter" action to achieve this target.
(6) Selecting a value: This action allows the agent to choose a specific value from a drop-down menu. We use the "select" action to achieve this target.
(7) End of task: Agents can signal the completion of the task or respond (e.g., an answer, task completion status, or indication of task impossibility). We use the "answer" action to achieve this target.
Figure 6: The actions in our action space.
Our datasets use the actions displayed in black color. The actions displayed in gray are not utilized, because their functions can be replaced by other actions in the same group. For example, we substitute the "type" and "paste" actions with the "input" action. Also, we replace the "drag" action (scroll bar dragging) with the "scroll" action. To facilitate the logging of information, we employ two-step actions (select_text and copy) rather than directly writing the texts. This approach is adopted to mitigate extensive time costs associated with crowd-sourcing annotation.
Actions Conversation from AITW dataset.
During the contribution of our GUIAct-smartphone dataset, we use a subset of the AITW [31] dataset, a large smartphone navigation data with General/Install/Google-Apps/Single/WebShopping tags. We use a subset of the partition with the "General" tag. To keep the consistency of our action space, we convert their actions: (1) dual-point gesture: we split the "dual-point gesture" actions into "tap" and "swipe" by the distance of the touch point and the left point. Then, we use the touch point as the "point" parameter of the "tap" action, and use both points as the "dual-point" parameter of the "swipe" action. (2) type: we rename the "type" action as the "input" action, and reserve the "text" parameter of "type" as the "text" parameter of "input". (3) enter: we reserve the "enter" action. (4) go back/home: we reflect on these two actions to "tap" actions. We find the position of the two buttons on the bottom navigation bar as the "point" parameter of the "tap“ action. (5) task complete/impossible: we convert these two actions to "answer" actions, and the "task complete" and "task impossible" are the "text" parameters of the "answer" action.
The Format of Actions.
Two types of formats need to be considered for actions: parse format and position format. We need a structured text format to extract actions from GUI agents’ natural language responses, referred as to the parse format. Also, we need an absolute or related format to express the position information of the actions, referred as to the position format.
As for parse format, we try several formats such as JSON and YAML, as shown in Figure 7. In our experiments, we utilize JSON format as the action parse format for Qwen-VL and Fuyu-8B. For MiniCPM-V, we opt for the CSV format, which closely resembles natural language conventions.
Figure 7: The four different parse formats for action parsing.
As for position format, we use different position formats based on the models’ preferences: absolute position format for Fuyu-8B and related position format for Qwen-VL. We also have a variant of related position formats, used for MiniCPM-V, which is scaling the related value to [0,1000) as integers (e.g., "<box>481 565 506 592</box>"). We show the position formats and the four parameters in Table 6.
Table 6: The position and distance formats. An element’s position is composed of its left-top point and right-down point and can be expressed as <box>x1, y1, x2, y2</box>. A point’s position is composed of its x and y coordinates, expressed as <point>x, y</point>. The position of dual points is composed of its start point and end point. The scroll parameter’s distance information uses pixel values. There are "down" and "right" two parameters to signal y dim and x dim respectively. If the action is scrolling up, the value of the "down" parameter is negative.
Actions Param. Absolute Related
Click, Hover, Select element <box>657, 434, 691, 455</box> <box>0.481, 0.565, 0.506, 0.592</box>
Tap point <point>362, 1412</point> <point>0.503, 0.981</point>
Swipe, Select_Text dual points from: <point>884, 236</point> from: <point>0.730, 0.311</point>
to: <point>958, 233</point> to: <point>0.791, 0.307</point>
Scroll scroll down 496 right 0 down 0.65 right 0
A.2 LLM-Auto Annotation
We use LLMs to help our contribution processing for GUIAct (web-single), GUIAct (web-multi), and GUIChat datasets.
GUIAct (web-single): Website Selection.
We use GPT-4 to gather website scenarios and website URLs, the prompts are shown in Figure 8.
Figure 8: The prompts for acquiring website scenarios and website URLs.
GUIAct (web-single): LLM-Auto Annotation.
We use GPT-4V to obtain instruction-action pairs for each web screenshot. To help GPT-4V comprehend both the content and the index of the elements, we provide two images for each request. As shown in Figure 9, one image is the origin screenshot without any mask box, and the other image is a revised screenshot with an element index and mask box on each element. Due to that, GPT-4V can get full information on the website viewport in the first image and get the element ID information in the second image.
Figure 9: The prompts and images for acquiring question-action pairs by GPT-4V.
GUIAct (web-multi): Questions Acquisition.
We use LLMs (GPT-3.5 and Claude2) to generate English and Chinese questions for English and Chinese websites respectively. These questions all have special answers due to the information on the websites. The prompt template is shown in Figure 10.
Figure 10: The prompt template for acquiring questions about information inquiry on a designated website.
GUIChat: Question-Answer Pair Generation.
We use GPT-4 and employ a pipeline for generating both single-turn and multi-turn question-and-answer pairs. For single-turn question-and-answer generation, the process begins with generating 4 types of questions, followed by generating answers and answers with grounding. As for multi-turn question-and-answer generation, it involves generating one round of question and answer per turn based on the history of the conversation, with a maximum of five rounds. The prompts are shown in Figure 11.
Figure 11: The prompts for generating the GUIChat dataset.
A.3 Crowd-sourced Annotation
Instructions Given to Participants.
We have a data annotation handbook for the annotation task for GUIAct (web-multi) data. We show the main content in Figure 12.
Figure 12: The main content of our data annotation handbook for GUIAct (web-multi).
Cost.
(1) We pay 0.6 RMB per item to check the LLM automatic annotation samples in the GUIAct (web-single) data. The checking efficiency of each item is about 40 seconds. So, the hourly salary we pay the annotator is 54 RMB. (2) We pay 5 RMB per item to annotate GUIAct (web-multi) data, and the annotation efficiency is approximately 5 minutes per item. So, the hourly salary we pay the annotator is 60 RMB. The prices we offer are higher than the minimum hourly wage in Beijing, China (13.91 RMB).
Appendix B Experiments
B.1 Training Details
Dataset Partitions.
Our GUICourse includes three datasets: GUIEnv, GUIAct, and GUIChat. (1) GUIEnv has 10M samples for pre-training (GUIEnv-global) and 0.7M samples for SFT (GUIEnv-local). Meanwhile, we split 1.8k samples in GUIEnv-local as test data for OCR and grounding abilities evaluation. (2) GUIAct is composed of three partitions: GUIAct (web-single), GUIAct (web-multi), and GUIAct (smartphone). GUIAct (web-single) is a dataset with single-step actions in website scenarios, which has about 67k instructions automatically generated by 13k screenshots. GUIAct (web-multi) is a dataset with multi-step actions and 5,696 complex instructions, which are annotated by humans. Also, we simply 9,157 instructions from AITW for smartphone scenarios, referred to as GUIAct (smartphone). There are a total of 67k, 44k, and 67k samples for the three partitions, and we split about 1.4k, 1k, and 2k samples as testing data. The residual samples are used for training. (3) GUIChat has 44k single-turn QA pairs and 6k multi-turn dialogues. We use all the samples for SFT.
Training Stages.
Our training process can be split into two stages due to the training targets. In the training Stage1, we mainly improve the OCR and text-grounding abilities of VLMs, especially in complex website scenarios. In the training Stage2, we mainly enhance the VLMs’ knowledge about GUI systems, such as the function of elements and control actions.
The whole training pipeline is fixed, but the training details are tailor-made for different VLMs, considering their uneven fundamental abilities. For Qwen-VL, which has strong fundamental abilities, we only execute the second stage of training for it. For VLMs that have weak text-grounding abilities (e.g., Fuyu-8B) and have no text-grounding ability (e.g., MiniCPM-V), we use supervised fine-tuning (using GUIEnv-local data) and pretraining (using GUIEnv-global data) methods respectively. Meanwhile, we will adjust our training stage depending on actual training objectives. For navigation-only agents, we only use the GUIAct data in the second training stage. We will mix the GUIChat data for the external functions such as the multi-turns conversation, etc.
Resources Cost.
For each agent, We use 1 $\times$ GPU A100 for inference, 8 $\times$ GPU A100 for SFT, and 128 $\times$ GPU A100 for pre-training. As for MiniCPM-V (3.1B), it takes about 2 days for the whole pre-training stage using the GUIEnv dataset and about 8 hours for the SFT stage using GUIAct and GUIChat datasets with 3 epochs. Moreover, it takes several hours for the evaluation. As for Qwen-VL (9.6B) and Fuyu-8B (9.4B), we only do SFT and inference processes. It takes about 1 day for training with the same SFT data.
B.2 The Evaluation of Actions.
Type Exact Match Score.
We use Type Exact Match Score (Type EM) to calculate the accuracy for predicting the names of actions, and not care about the parameters of the actions. The Type EM score is also called action type match [48] or intend match [22] in other works.
Table 7: The metrics of different actions.
Action Param. Examples Score Is Success
Click element <box>32, 20, 56, 44</box> IoU EM(el.)
Tap point <point>183, 505</point> distance distance < 0.14
Hover element <box>299, 118, 347, 166</box> IoU EM(el.)
Input text "French butter cake" F1 F1 > 0.5
Scroll pixel down 500 right 0 direction direction
Swipe dual points from <point1> to <point2> direction direction
Select Text dual points from <point1> to <point2> IoU IoU > 0.5
Copy - - EM(type) EM(type)
Enter - - EM(type) EM(type)
Select element, text the <element> value to "Female" IoU + F1 EM(el.) + F1
Answer text "On April 16, 2021." F1 F1 > 0.5
Figure 13: The IoU score and the element attach method. The left partition is the metrics of IoU scores and the right partition is the sketch of the element attach method. We first calculate the center point of the predicted element, and then calculate the distances (the gray dashed line) from this point to all the candidate elements. We select the element with the minimum distance as the final predicted element (the element "C" in this figure).
Action Score and Success Rate.
We calculate action scores and success rates for every action, as shown in Figure 7. The action scores are calculated similarity between the predicted actions and the golden actions. However, the action success rates judge if the predicted actions achieve the same function as the golden actions. We design different rules to calculate the action scores and success rates for various actions in our action space.
(1) "click" and "hover": For evaluating the two actions, we employ the Intersection over Union (IoU) score of their respective element regions (referred to as boxes) as their action scores. To determine the action success rate, we compare the predicted elements’ IDs with the golden elements’ IDs using an exact match criterion. It is worth noting that we associate the predicted region with the element having the minimum distance, as illustrated in Figure 13. Generally, the success rates of "click" and "hover" actions (referred to as Cli.Acc) tend to be higher than their corresponding action scores.
(2) "tap": For the "tap" action, there are no candidate or golden regions available in the datasets. Instead, we determine the success of the action by calculating the distance between the predicted point and the golden point using the related position format. If the distance is less than 0.14, the action is considered successful, and the action score is calculated as $1-\frac{distance}{0.14}$. Otherwise, the action score and success rate are set to $0.0$.
(3) "input" and "answer": We utilize the F1 score as the action score. If the action score exceeds $0.5$, we consider the action to be successful.
(4) "scroll" and "swipe": We convert the parameters of these actions into four directions: up, down, right, and left. If the predicted and golden results have consistent directions, the action is deemed successful, and the action score is set to $1.0$. Otherwise, the action score and success rate are set to $0.0$.
(5) "copy" and "enter": There are no parameters in both actions. The action scores and the success rates are equal to the Type EM scores.
(6) “select_text”: This action involves selecting text within a swiped region. We evaluate the action score using the IoU metric. If the IoU score exceeds $0.5$, we consider the action to be successful.
(7) “select”: there are two parameters (element and text) in this action. We use the IoU and F1 score for its element parameter and text parameter respectively. We calculate the average of the two scores as the action score. The action is considered successful only if both parameters meet the success criteria.
Step Success Rate.
In our approach, we encounter scenarios where multiple actions need to be performed in a single step. For instance, based on a given screenshot, the required actions could be a combination of "click," "input," and "enter." Therefore, our datasets necessitate models that can predict one or multiple actions in a single step. To Evaluate the performance of a single step, (1) we first compare the top-n predicted actions with the golden actions, where $n$ represents the number of golden actions. If the number of predicted actions is less than $n$, the scores of missing actions are set to $0.0$. (2) Then, we calculate the average action score for each action that is being compared, providing an overall action score. (3) Finally, we calculate the step Success Rate (StepSR). We consider a step successful only if all the actions within that step are successfully executed.
B.3 Case Study
We show a long-chain example executing in the simulated smartphone environment in Figure 14.
Figure 14: An example of our GUI agent executing the instruction with multiple steps in the simulated smartphone environment.
Appendix C Limitations, Ethical Considerations, and Societal Impacts
Limitations.
We use pretraining and supervised fine-tuning for training GUI agents using various VLMs. However, these might not be enough to achieve assistant-level GUI agents. We are considering the reinforcement learning methods such as RLHF in the future. We provide training data about website environments and fuse data about smartphone environments in our training. However, to build a general GUIAgent, only website and smartphone scenarios are not enough. We would like to build or gather more GUI instruction data about computer systems and professional software in the future.
Societal Impacts.
The screenshots in our GUIEnv dataset are collected from C4, we can’t guarantee there is no offensive content in these website screenshots.
Ethical Considerations.
We provide fair wages for the annotators when constructing our datasets, and there is no personally identifiable information. When using existing assets, we cite the creators and use the open-sourced datasets or models.